The perimeter of a rectangle is 26 ft. The length is 3 ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width.
Answer by stanbon(75887)

The perimeter of a rectangle is 26 ft. The length is 3 ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system.
--------------------
2(L + W) = 26
L = W + 3
------
Rearrange:
2W+3 = 13
Width = 5 ft.
------
Length = 8 ft
=======================
Cheers,
Stan H.
=========================